OpenStudy (anonymous):

What is the solution of the system? y=-3x 3x+2y=6 A. (2,6) B. (-1,3) C. (-2,6) D. (1.5, -4.5)

OpenStudy (anonymous):

you have got to be kidding...

OpenStudy (solomonzelman):

y=-3x 3x+2y=6 3x+2(-3x)=6 3x-6x=6 -3x=6 x=-2 can you solve for y?

OpenStudy (solomonzelman):

Hint: plug in the x value that you got, into the first equation.

OpenStudy (anonymous):

Is y 6?

OpenStudy (solomonzelman):

Good job!

OpenStudy (solomonzelman):

so if x=-2 y=6 and the discordant is (x,y) your's would be (-2,6) MAKES SENSE?

OpenStudy (solomonzelman):

I meant coordinate, not discordant, sorry.

OpenStudy (anonymous):

Yes! Thank you!

OpenStudy (solomonzelman):

Anytime!